# DiFuse-Net: RGB and Dual-Pixel Depth Estimation using Window Bi-directional Parallax Attention and Cross-modal Transfer Learning

### v1

@@ I INTRODUCTION
Additionally, since there is no RGB-DP-D dataset with high-quality ground-truth depth, we present a new method to overcome this challenge using a symmetric stereo setup with two smartphones. Acknowledging the non-rigid nature of smartphone lens systems, we employ regular calibration and rectification protocol for each capture session. We then train an AI stereo disparity estimation method on synthetic data to generate high-quality ground-truth depth maps, ensuring the method is robust to minor rectification errors ($<3$ pixels) inevitable in our setup. Our scalable dataset capture setup and method leads to $5000$ novel training samples, and $700$ novel test samples, which is one of its kind in the literature. Fig. 2 shows samples from our new dataset named Dual-Camera Dual-Pixel (DCDP) dataset.
@@
    A new and scalable method to obtain RGB-DP-D dataset with high quality depth ground-truth using a custom symmetric stereo camera setup made with two smartphone devices. The ground-truth depth maps in our dataset are significantly dense and have better quality than the existing dataset in the literature [11]. Our new dataset named DCDP, containing $5000$ novel training samples, and $700$ novel test samples, will support future research on RGB-DP based depth estimation.
Figure 1: DP disparities from a sample captured with a Google Pixel $3$ smartphone. Note: (1). We refer to the DP images as ‘left’ and ‘right’ for consistency with prior DP literature, which often utilizes DSLR cameras, although the physical disparity in most smartphones is vertical due to the sensor design and arrangement. (2). The grayscale representation of DP images reflects the limitation of DP sensing to the green channel of the Bayer sensor in smartphones for cost savings.
@@ II RELATED WORK
Garg et al. [11] were the first to propose an RGB-DP based depth estimation method. The authors designed a lightweight network architecture named DPNet, which takes the channel-wise concatenated RGB and DP images as input and is trained end-to-end using an affine invariant loss function. To acquire the training dataset, they created a custom camera rig using five Google Pixel $3$ smartphones, and the ground-truth depth map was computed using multi-view stereo techniques. The ground-truth depth maps in their dataset are sparse and often erroneous, as shown in Fig. 2. Therefore, it limits the accuracy and quality of the depth estimation. Moreover, the authors do not focus on specialized processing of the DP images to compute the defocus disparity information, which makes it less effective.
Zhang et al. [17] extended the method in [11] by additionally using dual cameras. They kept the baseline of the dual cameras orthogonal to the DP baseline to obtain complementary disparity information from the two sources for more accurate depth estimation. Pan et al. [18] proposed a method for DP image simulation for DSLR images and trained a multi-task model to estimate depth and perform image deblurring simultaneously. In [19], the authors proposed a method for modeling defocus disparity in DSLR DP images, operating on patches and stitching the outputs to recover the disparity map. The authors also provide $100$ RGB-DP-Disparity triplets of simple lab scenes using a DSLR, whereas some handful RGB-DP images from Pixel smartphone were used only for qualitative evaluation. As it is pointed in [19], it is important to note that smartphone and DSLR DPs are significantly different, therefore, model trained on DSLR dataset cannot be evaluated on smartphone dataset and vice-versa.
@@ III-A RGB Encoder
The RGB encoder processes the input images $I\in\mathcal{R}^{H\times W\times 3}$. It employs an EfficientNet-Lite3333 backbone [21], initialized with pre-trained ImageNet [22] weights, for feature extraction. The network progressively downsamples the image by a factor of $32$ using convolutional blocks. Each block halves the resolution and increases the number of channels while utilizing inverted residual blocks [23] for efficient computation. We incorporate an additional inverted residual block for further downsampling to $1/64$ of the original size, enlarging the receptive field and enhancing the global contextual information.
@@ III-B DP Encoder
The DP encoder is designed to extract defocus disparity cues from DP image pairs. Exploiting the fact that DP images exhibit disparities along the epipolar line, we employ modified parallax attention [24] oriented along this axis to capture these cues effectively. Due to the limited disparity range (typically $-8$ to $+8$ pixels) inherent to smartphone cameras with small apertures, correspondence matching is constrained to a local neighborhood.
@@ IV DUAL-CAMERA DUAL-PIXEL DATASET GENERATION
The Google Pixel $2/3$ dataset [11], while valuable, suffers from sparse and often erroneous ground-truth depth (Fig. 2). To address this, we propose a methodology for constructing a high-quality RGB-DP-D dataset. Existing depth sensors [9, 1] lack DP capture and pose alignment challenges. Unlike the complex multi-camera setup in [11], we employ a simplified symmetric stereo configuration using two DP sensor enabled smartphones, leveraging recent AI stereo disparity estimation advancements [4, 5]. Our setup prioritizes two key criteria for high-quality data:
@@ IV-A Symmetric Stereo Camera Setup
Symmetric stereo cameras simplify calibration and rectification compared to wide/telephoto combinations [8]. We employed two Samsung Galaxy S23232323 Ultra smartphones in a symmetric stereo configuration (Fig. 5), using the front cameras to minimize baseline ($2.5cm$) and occlusions. Rigid camera holders and a metal support ensured stability. Simultaneous capture was achieved using Galaxy S-pens and USB-C camera switch.
@@ IV-B Stereo Rectification
Our camera setup primarily exhibits horizontal disparity. To ensure precise alignment, we perform stereo calibration and rectification before each capture session, utilizing $30$-$40$ checkerboard samples. Each session yields $120$-$150$ images (Fig. 6). Such a routine protocol mitigates potential positional shifts in long capture sessions. Calibration and rectification are performed at half-resolution, with the resulting coordinate mapping applied to full-resolution images. Rectified stereo pairs are processed through an AI stereo disparity estimation pipeline.
To address the misalignment between the estimated disparity map (in the rectified plane) and the captured DP images, we project the disparity map back to the original plane, thus preserving DP cues. This is preferable to rectifying DP images, which would distort DP cues. Although reversing distortion is less precise than inverting pose parameters, our minimal distortion coefficients allow us to approximate reverse distortion by inverting their signs. A $40$-pixel border crop mitigates any residual inaccuracies (Fig. 7).
@@
Figure 7: (a) Stereo pairs. (b) Rectified stereo pairs. (c) Estimated disparity map in the rectified plane. (d) Disparity map projected back to the original plane along with $40$-pixel border crop (the same crop is later applied to original RGB-DP images for alignment).
@@ IV-C Ground-truth Depth Generation
We leverage advances in AI stereo disparity estimation [4, 5] to generate accurate disparity maps from rectified stereo pairs. Our model, trained on a large synthetic dataset [27] augmented with slight vertical distortions, ensures robustness to minor rectification errors. Back-projecting the estimated disparity map and applying a $40$-pixel border crop yields aligned RGB-DP-D samples with preserved DP cues.
@@
Finally, Fig. 9 provides a qualitative assessment of the ground-truth depth accuracy in the DCDP dataset. By rendering the point clouds from a novel viewpoint, we highlight the geometric consistency and absence of artifacts in the reconstructed $3$D structure.
@@
Figure 9: Qualitative assessment of the ground-truth depth accuracy in the DCDP dataset. Point clouds are rendered from a novel viewpoint to highlight the geometric consistency in the reconstructed $3$D structure.
@@ V-A Datasets
Google DP Dataset: Google DP dataset [11] consists of $12,530$ training and $684$ test RGB-DP-D samples. Since each scene is captured by $5$ cameras in the multi-camera rig, effectively there are $2506$ novel images in this dataset.
DCDP: Our DCDP dataset consists of $5000$ novel training and $700$ novel test RGB-DP-D samples.
@@ V-B Implementation
We used the PyTorch [28] framework for implementation. For training, we used Adam optimizer [29] with momentum of $0.9$. We employed a polynomial learning rate decay scheduler with power term set to $0.9$, and an initial learning rate of $1\times 10^{-4}$. Similar to the loss function in [13], we combine the scale-invariant mean absolute error with a scale-invariant gradient matching term. The gradient matching term helps to preserve sharp edges and boundaries [13]. Also, similar to [11] the model is trained to predict the inverse depth.
@@
In Eq. 3, $d$ signifies the ground-truth inverse depth, whereas $\hat{d}$ denotes the predicted affine invariant depth. $\lambda$ is a hyper-parameter, set to $30$ across all experiments. MAE($\cdot$) and Grad($\cdot$) represent the mean absolute error and the gradient loss functions, respectively.
@@ V-C Evaluation Metrics
Similar to [11], we use Spearman’s Rank Correlation Coefficient (SRCC) to evaluate the ordinal correctness of the estimated depth map with ground-truth depth, and affine-invariant versions of mean absolute error (MAE) and root mean squared error (RMSE), denoted by AIWE $1$ and AIWE $2$ respectively.
@@ VI-A Evaluation and Comparison
Tab. I shows detailed quantitative comparison of DiFuse-Net with other baselines, viz., DPNet [11], Baseline, and Stereo Baseline [30]. We report the latest DPNet metrics provided on their official GitHub page due to their release of modified train and test datasets. This ensures a fair evaluation of our method against DPNet. Baseline is the enhanced DPNet model by increasing the convolutional layers, aligning the total number of parameters with that of DiFuse-Net, to ensure a more fair comparison. Stereo Baseline replaces WBiPAM with a traditional stereo matching cost-volume approach [30, 5, 8], providing a comparison with established stereo methods. Total number of parameters of Stereo Baseline are aligned with that of DiFuse-Net to ensure a fair comparison.
@@ VI-B Ablation Study
DPNet[11] $0.1520$ $0.0181$ $0.0268$
Baseline $0.0927$ $0.0142$ $0.0218$
Stereo Baseline [30] $0.0911$ $0.0137$ $0.0216$
@@
WBiPAM No WBiPAM $0.0865$ $0.0132$ $0.0210$
No Window WBiPAM $0.0912$ $0.0137$ $0.0214$
Unidirectional WBiPAM $0.1454$ $0.0191$ $0.0276$
Fusion Module Pixel-wise Fusion $0.0919$ $0.0138$ $0.0214$
Channel-wise Fusion $0.0855$ $0.0129$ $0.0204$
DP Encoder 5 Layer DP Encoder $0.0950$ $0.0138$ $0.0215$
4 Layer DP Encoder $0.0889$ $0.0135$ $0.0210$
3 Layer DP Encoder $0.0868$ $0.0134$ $0.0210$
1 Layer DP Encoder $0.0859$ $0.0135$ $0.0210$
CmTL DiFuse-Net w/o CmTL $0.0833$ $0.0129$ $0.0205$
@@
MiDaS v3.13.13.13.1 $\text{BEiT}_{\text{L}-512}$ [13] $0.0971$ $0.0168$ $0.0267$
ZoeDepth [31] $0.1168$ $0.0272$ $0.0379$
@@ VI-C Additional Evaluation
We conducted a comparative analysis between DiFuse-Net and monocular depth estimation models trained on large-scale RGB-D datasets, viz., MiDaS v3.13.13.13.1 $\text{BEiT}_{\text{L}-512}$ [13] which comprises of $345$M parameters, and ZoeDepth [31]. ZoeDepth uses MiDaS as backbone, and uses a metric depth prediction head. Best numbers were obtained by the ZoeD-M12121212-N (trained on NYU V2222) variant. Our analysis revealed that despite DiFuse-Net having substantially fewer paramaters ($9.9$M), its depth map quality was better than that of MiDaS, owing to the effective use of DP images in an improved model architecture. Table  III shows the quantitative score comparison of DiFuse-Net with MiDaS, and ZoeDepth indicating better scores for our model. The qualitative comparisons are presented in  11. where our model exhibits enhanced details.
@@ VI-D Evaluation on the DCDP Dataset
DPNet [11] $0.1522$ $0.0087$ $0.0128$
Baseline $0.0928$ $0.0068$ $0.0104$
Stereo Baseline [30] $0.0912$ $0.0066$ $0.0098$

### v2

@@ I INTRODUCTION
Additionally, since there is no RGB-DP-D dataset with high-quality ground-truth depth, we present a new method to overcome this challenge using a symmetric stereo setup with two smartphones. Acknowledging the non-rigid nature of smartphone lens systems, we employ regular calibration and rectification protocol for each capture session. We then train an AI stereo disparity estimation method on synthetic data to generate high-quality ground-truth depth maps, ensuring the method is robust to minor rectification errors (<3<3< 3 pixels) inevitable in our setup. Our scalable dataset capture setup and method leads to 500050005000 novel training samples, and 700700700 novel test samples, which is one of its kind in the literature. Fig. 2 shows samples from our new dataset named Dual-Camera Dual-Pixel (DCDP) dataset.
@@
    A new and scalable method to obtain RGB-DP-D dataset with high quality depth ground-truth using a custom symmetric stereo camera setup made with two smartphone devices. The ground-truth depth maps in our dataset are significantly dense and have better quality than the existing dataset in the literature [11]. Our new dataset named DCDP, containing 500050005000 novel training samples, and 700700700 novel test samples, will support future research on RGB-DP based depth estimation.
Figure 1: DP disparities from a sample captured with a Google Pixel 333 smartphone. Note: (1). We refer to the DP images as ‘left’ and ‘right’ for consistency with prior DP literature, which often utilizes DSLR cameras, although the physical disparity in most smartphones is vertical due to the sensor design and arrangement. (2). The grayscale representation of DP images reflects the limitation of DP sensing to the green channel of the Bayer sensor in smartphones for cost savings.
@@ II RELATED WORK
Garg et al. [11] were the first to propose an RGB-DP based depth estimation method. The authors designed a lightweight network architecture named DPNet, which takes the channel-wise concatenated RGB and DP images as input and is trained end-to-end using an affine invariant loss function. To acquire the training dataset, they created a custom camera rig using five Google Pixel 333 smartphones, and the ground-truth depth map was computed using multi-view stereo techniques. The ground-truth depth maps in their dataset are sparse and often erroneous, as shown in Fig. 2. Therefore, it limits the accuracy and quality of the depth estimation. Moreover, the authors do not focus on specialized processing of the DP images to compute the defocus disparity information, which makes it less effective.
Zhang et al. [17] extended the method in [11] by additionally using dual cameras. They kept the baseline of the dual cameras orthogonal to the DP baseline to obtain complementary disparity information from the two sources for more accurate depth estimation. Pan et al. [18] proposed a method for DP image simulation for DSLR images and trained a multi-task model to estimate depth and perform image deblurring simultaneously. In [19], the authors proposed a method for modeling defocus disparity in DSLR DP images, operating on patches and stitching the outputs to recover the disparity map. The authors also provide 100100100 RGB-DP-Disparity triplets of simple lab scenes using a DSLR, whereas some handful RGB-DP images from Pixel smartphone were used only for qualitative evaluation. As it is pointed in [19], it is important to note that smartphone and DSLR DPs are significantly different, therefore, model trained on DSLR dataset cannot be evaluated on smartphone dataset and vice-versa.
@@ III-A RGB Encoder
The RGB encoder processes the input images $I\in\mathcal{R}^{H\times W\times 3}$. It employs an EfficientNet-Lite333 backbone [21], initialized with pre-trained ImageNet [22] weights, for feature extraction. The network progressively downsamples the image by a factor of 323232 using convolutional blocks. Each block halves the resolution and increases the number of channels while utilizing inverted residual blocks [23] for efficient computation. We incorporate an additional inverted residual block for further downsampling to 1/641/641 / 64 of the original size, enlarging the receptive field and enhancing the global contextual information.
@@ III-B DP Encoder
The DP encoder is designed to extract defocus disparity cues from DP image pairs. Exploiting the fact that DP images exhibit disparities along the epipolar line, we employ modified parallax attention [24] oriented along this axis to capture these cues effectively. Due to the limited disparity range (typically 8-8- 8 to +8+8+ 8 pixels) inherent to smartphone cameras with small apertures, correspondence matching is constrained to a local neighborhood.
@@ IV DUAL-CAMERA DUAL-PIXEL DATASET GENERATION
The Google Pixel 2/32/32 / 3 dataset [11], while valuable, suffers from sparse and often erroneous ground-truth depth (Fig. 2). To address this, we propose a methodology for constructing a high-quality RGB-DP-D dataset. Existing depth sensors [9, 1] lack DP capture and pose alignment challenges. Unlike the complex multi-camera setup in [11], we employ a simplified symmetric stereo configuration using two DP sensor enabled smartphones, leveraging recent AI stereo disparity estimation advancements [4, 5]. Our setup prioritizes two key criteria for high-quality data:
@@ IV-A Symmetric Stereo Camera Setup
Symmetric stereo cameras simplify calibration and rectification compared to wide/telephoto combinations [8]. We employed two Samsung Galaxy S232323 Ultra smartphones in a symmetric stereo configuration (Fig. 5), using the front cameras to minimize baseline ($2.5cm$) and occlusions. Rigid camera holders and a metal support ensured stability. Simultaneous capture was achieved using Galaxy S-pens and USB-C camera switch.
@@ IV-B Stereo Rectification
Our camera setup primarily exhibits horizontal disparity. To ensure precise alignment, we perform stereo calibration and rectification before each capture session, utilizing 303030-404040 checkerboard samples. Each session yields 120120120-150150150 images (Fig. 6). Such a routine protocol mitigates potential positional shifts in long capture sessions. Calibration and rectification are performed at half-resolution, with the resulting coordinate mapping applied to full-resolution images. Rectified stereo pairs are processed through an AI stereo disparity estimation pipeline.
To address the misalignment between the estimated disparity map (in the rectified plane) and the captured DP images, we project the disparity map back to the original plane, thus preserving DP cues. This is preferable to rectifying DP images, which would distort DP cues. Although reversing distortion is less precise than inverting pose parameters, our minimal distortion coefficients allow us to approximate reverse distortion by inverting their signs. A 404040-pixel border crop mitigates any residual inaccuracies (Fig. 7).
@@
Figure 7: (a) Stereo pairs. (b) Rectified stereo pairs. (c) Estimated disparity map in the rectified plane. (d) Disparity map projected back to the original plane along with 404040-pixel border crop (the same crop is later applied to original RGB-DP images for alignment).
@@ IV-C Ground-truth Depth Generation
We leverage advances in AI stereo disparity estimation [4, 5] to generate accurate disparity maps from rectified stereo pairs. Our model, trained on a large synthetic dataset [27] augmented with slight vertical distortions, ensures robustness to minor rectification errors. Back-projecting the estimated disparity map and applying a 404040-pixel border crop yields aligned RGB-DP-D samples with preserved DP cues.
@@
Finally, Fig. 9 provides a qualitative assessment of the ground-truth depth accuracy in the DCDP dataset. By rendering the point clouds from a novel viewpoint, we highlight the geometric consistency and absence of artifacts in the reconstructed  [^3] D structure.
@@
Figure 9: Qualitative assessment of the ground-truth depth accuracy in the DCDP dataset. Point clouds are rendered from a novel viewpoint to highlight the geometric consistency in the reconstructed  [^3] D structure.
@@ V-A Datasets
Google DP Dataset: Google DP dataset [11] consists of 12,53012,53012 , 530 training and 684684684 test RGB-DP-D samples. Since each scene is captured by 555 cameras in the multi-camera rig, effectively there are 250625062506 novel images in this dataset.
DCDP: Our DCDP dataset consists of 500050005000 novel training and 700700700 novel test RGB-DP-D samples.
@@ V-B Implementation
We used the PyTorch [28] framework for implementation. For training, we used Adam optimizer [29] with momentum of 0.90.90.9. We employed a polynomial learning rate decay scheduler with power term set to 0.90.90.9, and an initial learning rate of $1\times 10^{-4}$. Similar to the loss function in [13], we combine the scale-invariant mean absolute error with a scale-invariant gradient matching term. The gradient matching term helps to preserve sharp edges and boundaries [13]. Also, similar to [11] the model is trained to predict the inverse depth.
@@
In Eq. 3, $d$ signifies the ground-truth inverse depth, whereas $\hat{d}$ denotes the predicted affine invariant depth. $\lambda$ is a hyper-parameter, set to 303030 across all experiments. MAE($\cdot$) and Grad($\cdot$) represent the mean absolute error and the gradient loss functions, respectively.
@@ V-C Evaluation Metrics
Similar to [11], we use Spearman’s Rank Correlation Coefficient (SRCC) to evaluate the ordinal correctness of the estimated depth map with ground-truth depth, and affine-invariant versions of mean absolute error (MAE) and root mean squared error (RMSE), denoted by AIWE 111 and AIWE 222 respectively.
@@ VI-A Evaluation and Comparison
Tab. I shows detailed quantitative comparison of DiFuse-Net with other baselines, viz., DPNet [11], Baseline, and Stereo Baseline [30]. We report the latest DPNet metrics provided on their official GitHub page due to their release of modified train and test datasets. This ensures a fair evaluation of our method against DPNet. Baseline is the enhanced DPNet model by increasing the convolutional layers, aligning the total number of parameters with that of DiFuse-Net, to ensure a more fair comparison. Stereo Baseline replaces WBiPAM with a traditional stereo matching cost-volume approach [30, 5, 8], providing a comparison with established stereo methods. More specifically, we use [30]. The total number of parameters of Stereo Baseline are aligned with that of DiFuse-Net to ensure a fair comparison.
@@ VI-B Ablation Study
DPNet[11] 0.15200.15200.1520 0.01810.01810.0181 0.02680.02680.0268
Baseline 0.09270.09270.0927 0.01420.01420.0142 0.02180.02180.0218
Stereo Baseline [30] 0.09110.09110.0911 0.01370.01370.0137 0.02160.02160.0216
@@
WBiPAM No WBiPAM 0.08650.08650.0865 0.01320.01320.0132 0.02100.02100.0210
No Window WBiPAM 0.09120.09120.0912 0.01370.01370.0137 0.02140.02140.0214
Unidirectional WBiPAM 0.14540.14540.1454 0.01910.01910.0191 0.02760.02760.0276
Fusion Module Pixel-wise Fusion 0.09190.09190.0919 0.01380.01380.0138 0.02140.02140.0214
Channel-wise Fusion 0.08550.08550.0855 0.01290.01290.0129 0.02040.02040.0204
DP Encoder 5 Layer DP Encoder 0.09500.09500.0950 0.01380.01380.0138 0.02150.02150.0215
4 Layer DP Encoder 0.08890.08890.0889 0.01350.01350.0135 0.02100.02100.0210
3 Layer DP Encoder 0.08680.08680.0868 0.01340.01340.0134 0.02100.02100.0210
1 Layer DP Encoder 0.08590.08590.0859 0.01350.01350.0135 0.02100.02100.0210
CmTL DiFuse-Net w/o CmTL 0.08330.08330.0833 0.01290.01290.0129 0.02050.02050.0205
@@
MiDaS v3.13.13.1 $\text{BEiT}_{\text{L}-512}$ [13] 0.09710.09710.0971 0.01680.01680.0168 0.02670.02670.0267
ZoeDepth [31] 0.11680.11680.1168 0.02720.02720.0272 0.03790.03790.0379
@@ VI-C Additional Evaluation
We conducted a comparative analysis between DiFuse-Net and monocular depth estimation models trained on large-scale RGB-D datasets, viz., MiDaS v3.13.13.1 $\text{BEiT}_{\text{L}-512}$ [13] which comprises of 345345345M parameters, and ZoeDepth [31]. ZoeDepth uses MiDaS as backbone, and uses a metric depth prediction head. Best numbers were obtained by the ZoeD-M121212-N (trained on NYU V222) variant. Our analysis revealed that despite DiFuse-Net having substantially fewer paramaters (9.99.99.9M), its depth map quality was better than that of MiDaS, owing to the effective use of DP images in an improved model architecture. Table  III shows the quantitative score comparison of DiFuse-Net with MiDaS, and ZoeDepth indicating better scores for our model. The qualitative comparisons are presented in  11. where our model exhibits enhanced details.
@@ VI-D Evaluation on the DCDP Dataset
DPNet [11] 0.15220.15220.1522 0.00870.00870.0087 0.01280.01280.0128
Baseline 0.09280.09280.0928 0.00680.00680.0068 0.01040.01040.0104
Stereo Baseline [30] 0.09120.09120.0912 0.00660.00660.0066 0.00980.00980.0098